O 2003 Doney /Northampton Community Colle sics 151 ge e theoretical value of the axial magnetic field inside the coil using the measured current, length, and number of turns for the coil (for SE-8653 outer coil use N 2920 turns, length actual 11 Show your work here, including units and conversions. cm) Theoretical Magnetic Field Compare the theoretical value to the actual value using a percent difference. that could account for this percent difference? What are some factors 6.
Answer #1

For a solenoid, the axial magnetic field inside the coil is given by

B= \mu_{o} nI

where,

\mu_{o} = 4\pi \times 10^{-7}

n is number of turns of coils per unit length

n= \frac{N}{L}=\frac{2920}{0.11} turns /m

I is the measured current, which you have not mentioned in the question.

So, to calculate the value of magnetic field, please multiply as shown below, plugging in your measured value of current in place of I

B = \frac{4\times 3.14 \times 10^{-7} \times 2920 \times I}{0.11}

If you put the value of current in Amperes, you will get the value of magnetic field in Tesla, so the unit to the above obtained value will be T